25. January 2016: Monday

LDAP/IdM:

  • Updates:
    • No issues with last night’s update.
    • Issue from last week (in lookaside) resolved, merged.
  • Accounts:
    • No errors (0) in 50 new accounts.
    • Renamed FS account needed kerberos fix

Backups:

  • Save Groups:
    • The usual list
    • vcenter1: finally resolved the issue, the new network was not allowed to talk to the storage nodes. ┬áIt works now.
    • view-vc1: J. Greg believes he finally got that fixed.
  • Tapes:
    • Relabeled 60 more into Scratch pool
    • Reclaimed 3 for the vaults.
    • Fixed 7,684 SSIDs that cloning/staging have messed up the retention values on.
  • Problems:
    • “minor” network configuration issue on a core router isolated Tech Park for several minutes and caused split-brain issues and a backup to fail.
      • Updated /etc/hosts on the servers so they can all find each other should DNS become unavailable again.

Tags:

Leave a Reply

You must be logged in to post a comment.